OpenStudy (anonymous):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, and his speed decreases at a rate of 1.5 miles per hour. The equation gives the speed of the boat after David has rowed for x hours? can someone help y=4+1.5x y=4-1.5x y=6-1.5x

OpenStudy (jdoe0001):

ok, thanks for sharing that

OpenStudy (anonymous):

y=4+1.5x y=4-1.5x y=6-1.5x

OpenStudy (irishboy123):

GALILEO!!!!! is there a reference point? EINSTEIN!!!! which velocity is absolute?!?!